The area of a circle is 32pi square units.
Select one:
a. 8pi√2
b. Not enough information to find the circumference.
c. 2pi * 5.7
d. 35.5
e. 2pi√32

Answer by rothauserc(4718)

Area(A) of circle = pi*radius(r)^2
Circumference of circle (C) = pi*diameter(D)
we are given A = 32*pi, therefore
32*pi = pi*r^2
divide both sides of = by pi
r^2 = 32
r = square root of 32
now C = pi*D and D = 2r, therefore
C = 2*pi*square root of 32
answer is e